# Smooth Tests for Normality in ANOVA

### v3

@@ Appendix A Appendix
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1).$
@@
$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}.$
@@
$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left\{\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]\right$.
@@
$=$ $2a^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}$
$\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.28)
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right)-a_{1},\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)-a_{K}\right]^{\top}$
$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{W}_{j},$
@@
$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left\{\left(\pi_{k}(Z)-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2k}}{2}\left(e^{2}-1\right)\right\}$
$\quad\left\{\left(\pi_{l}(Z)-\mathrm{E}\left[\pi_{l}(Z)\right]\right)-d_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2l}}{2}\left(e^{2}-1\right)\right\}$
$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right),$
and the asymptotic distribution of $\tilde{\Psi}_{K}^{2}$ in (3.15) can be derived from (A.28) through similar arguments above. This completes the proof of Theorem 8.
The proof of Theorem 9 can be referred to that of Theorem 3 due to the similar arguments and techniques used. To obtain the results in Theorem 9, following the proof of Theorem 8, it suffices to show that under $H_{1L}$, $\mathrm{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, $\mathrm{E}[\pi_{k}(Z)\pi_{l}(Z)]=\delta_{kl}+o(1)$, $d_{1k}=c_{1k}+o(1)$, $d_{2k}=c_{2k}+o(1)$, $d_{3k}=c_{1k}+o(1)$, $d_{4k}=c_{2k}+o(1)$, $\mathrm{E}[e^{3}]=o(1)$, and $\mathrm{E}[e^{4}]=3+o(1)$, which can be found in the proof of Theorem 3. $\square$
Proof of Theorem 10.
We focus on the first case in Theorem 10, and the latter two cases can be verified analogously. We primarily follow well-established steps in the literature, for example, the proof of Theorem 2 in Ducharme and Lafaye de Micheaux, (2004), Lemma 1 in Kraus, (2007), and Propositions 6 and 7 in Duchesne et al., (2016). Note that $\{\hat{K}=k\}\subseteq\{N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\}$ for $k^{\prime}\neq k$, then we have
$\mathrm{P}\left(\hat{K}=k\right)\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\right),\quad k^{\prime}\neq k.$ (A.29)
Under $H_{0}$, in order to show $\mathrm{P}(\hat{K}=1)\to 1$ as $N\to\infty$, we first show the contrary such that $\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 0$. In fact, from (A.29) we have
$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)$
$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{1}-\log N\right)$
$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq-\log N\right)$
$=$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right).$
From Corollary 1 we know that under $H_{0}$, $N\hat{\Psi}^{2}_{k}\to_{d}\chi^{2}_{k}$ for each $k=1,\ldots,D$. Thus for $k=2,\ldots,D$, $\mathrm{P}(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N)\to 0$ as $N\to\infty$. Therefore,
$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)\leq\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right)\to 0,$
as $N\to\infty$, then $\mathrm{P}(\hat{K}=1)=1-\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 1$, and $N\hat{\Psi}^{2}_{\hat{K}}\to_{d}\chi^{2}_{1}$ follows.
@@
$\mathrm{P}\left(\hat{K}=k\right)$ $\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{K_{0}}-K_{0}\log N\right)$
$=\mathrm{P}\left(\hat{\Psi}^{2}_{k}-k\frac{\log N}{N}\geq\hat{\Psi}^{2}_{K_{0}}-K_{0}\frac{\log N}{N}\right).$
From Theorem 2 we know that as $N\to\infty$, $\hat{\Psi}^{2}_{k}\to_{p}0$ for $k=1,\ldots,K_{0}-1$, $\hat{\Psi}^{2}_{K_{0}}\to_{p}\bm{a}_{K_{0}}^{\top}\bm{\Sigma}_{K_{0}}^{-1}\bm{a}_{K_{0}}>0$ and $\log N/N\to 0$. Thus $\mathrm{P}(\hat{K}=k)\to 0$ for $k=1,\ldots,K_{0}-1$, and
$\lim_{N\to\infty}\mathrm{P}\left(\hat{K}\geq K_{0}\right)=1-\lim_{N\to\infty}\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)=1.$
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)$ $=\sum_{k=1}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x,\hat{K}=k\right)$
$\leq\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)+\sum_{k=K_{0}}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x\right)\to 0,$
since $\hat{\Psi}^{2}_{k}\to_{p}\bm{a}_{k}^{\top}\bm{\Sigma}_{k}^{-1}\bm{a}_{k}>0$ for $k=K_{0},\ldots,D$. Therefore, for any $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)=0,$
@@
Intuition behind the revised limiting null distribution (4.3).
We primarily follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), and Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case, as the latter two cases can be derived analogously. Note that
$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)=$ $\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)$
$+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
The third term on the right-hand side of (A), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ). The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
We first investigate the asymptotic distribution of $(N\hat{\Psi}^{2}_{1},N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1}))^{\top}$. The limiting distributions of $N\hat{\Psi}^{2}_{1},N\hat{\Psi}^{2}_{2}$ are functions of a bivariate normal vector $(R_{1},R_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\bm{\Sigma}_{2})$. We generally denote elements of $\bm{\Sigma}_{2}$ as $\begin{pmatrix}a&b\\ b&c\end{pmatrix}$ and $\rho=b/\sqrt{ac}$. The distribution of $(R_{1},R_{2})^{\top}$ can be constructed through two independent standard normal variables $G_{1},G_{2}$ if we denote $\tilde{R}_{1}=\sqrt{a}(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})$ and $\tilde{R}_{2}=\sqrt{c}G_{2}$, then $(\tilde{R}_{1},\tilde{R}_{2})^{\top}\overset{d}{=}(R_{1},R_{2})^{\top}$. Thus, $N\hat{\Psi}^{2}_{1}\to_{d}R_{1}^{2}/a\overset{d}{=}\tilde{R}_{1}^{2}/a=(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})^{2}$. For $N\hat{\Psi}_{2}^{2}$, straightforward computations yield that
$N\hat{\Psi}_{2}^{2}$ $\to_{d}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\bm{\Sigma}_{2}^{-1}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(c\tilde{R}_{1}^{2}-2b\tilde{R}_{1}\tilde{R}_{2}+a\tilde{R}_{2}^{2}\right)$
@@
Then $N(\hat{\Psi}_{2}^{2}-\hat{\Psi}_{1}^{2})\to_{d}(\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{2}$. Since $(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2},\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\mathbf{I}_{2})$, we obtain that $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ are both asymptotically $\chi_{1}^{2}$ distributed and are asymptotically independent.
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)\approx\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}>\log N\right)=0,$
by $N\hat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)$

### v4

@@ A.6 Proof of Theorem 8 and 9
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}(1).$
@@
$\widetilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)\overset{p}{\to}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}.$
@@
$\sqrt{N}\left(\widetilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left\{\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]\right$.
@@
$=$ $2a^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}$
$\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.28)
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right)-a_{1},\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)-a_{K}\right]^{\top}$
$\overset{d}{\to}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{W}_{j},$
@@
$\lambda_{kl}^{(j)}=$ $\mathbb{E}\left\{\left(\pi_{k}(Z)-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2k}}{2}\left(e^{2}-1\right)\right\}$
$\quad\left\{\left(\pi_{l}(Z)-\mathbb{E}\left[\pi_{l}(Z)\right]\right)-d_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2l}}{2}\left(e^{2}-1\right)\right\}$
$=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathbb{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right),$
and the asymptotic distribution of $\widetilde{\Psi}_{K}^{2}$ in (3.15) can be derived from (A.28) through similar arguments above. This completes the proof of Theorem 8.
The proof of Theorem 9 can be referred to that of Theorem 3 due to the similar arguments and techniques used. To obtain the results in Theorem 9, following the proof of Theorem 8, it suffices to show that under $H_{1L}$, $\mathbb{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, $\mathbb{E}[\pi_{k}(Z)\pi_{l}(Z)]=\delta_{kl}+o(1)$, $d_{1k}=c_{1k}+o(1)$, $d_{2k}=c_{2k}+o(1)$, $d_{3k}=c_{1k}+o(1)$, $d_{4k}=c_{2k}+o(1)$, $\mathbb{E}[e^{3}]=o(1)$, and $\mathbb{E}[e^{4}]=3+o(1)$, which can be found in the proof of Theorem 3. $\square$
A.7 Proof of Theorem 10
We focus on the first case in Theorem 10, and the latter two cases can be verified analogously. We primarily follow well-established steps in the literature, for example, the proof of Theorem 2 in ducharme2004goodness, Lemma 1 in kraus2007data, and Propositions 6 and 7 in duchesne2016estimating. Note that $\{\widehat{K}=k\}\subseteq\{N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\}$ for $k^{\prime}\neq k$, then we have
$\mathbb{P}\left(\widehat{K}=k\right)\leq\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\right),\quad k^{\prime}\neq k.$ (A.29)
Under $H_{0}$, in order to show $\mathbb{P}(\widehat{K}=1)\to 1$ as $N\to\infty$, we first show the contrary such that $\sum_{k=2}^{D}\mathbb{P}(\widehat{K}=k)\to 0$. In fact, from (A.29) we have
$\sum_{k=2}^{D}\mathbb{P}\left(\widehat{K}=k\right)$
$\leq$ $\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{1}-\log N\right)$
$\leq$ $\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq-\log N\right)$
$=$ $\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\geq(k-1)\log N\right).$
From Corollary 1 we know that under $H_{0}$, $N\widehat{\Psi}^{2}_{k}\overset{d}{\to}\chi^{2}_{k}$ for each $k=1,\ldots,D$. Thus for $k=2,\ldots,D$, $\mathbb{P}(N\widehat{\Psi}^{2}_{k}\geq(k-1)\log N)\to 0$ as $N\to\infty$. Therefore,
$\sum_{k=2}^{D}\mathbb{P}\left(\widehat{K}=k\right)\leq\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\geq(k-1)\log N\right)\to 0,$
as $N\to\infty$, then $\mathbb{P}(\widehat{K}=1)=1-\sum_{k=2}^{D}\mathbb{P}(\widehat{K}=k)\to 1$, and $N\widehat{\Psi}^{2}_{\widehat{K}}\overset{d}{\to}\chi^{2}_{1}$ follows.
@@ A.7 Proof of Theorem 10
$\mathbb{P}\left(\widehat{K}=k\right)$ $\leq\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{K_{0}}-K_{0}\log N\right)$
$=\mathbb{P}\left(\widehat{\Psi}^{2}_{k}-k\frac{\log N}{N}\geq\widehat{\Psi}^{2}_{K_{0}}-K_{0}\frac{\log N}{N}\right).$
From Theorem 2 we know that as $N\to\infty$, $\widehat{\Psi}^{2}_{k}\overset{p}{\to}0$ for $k=1,\ldots,K_{0}-1$, $\widehat{\Psi}^{2}_{K_{0}}\overset{p}{\to}\bm{a}_{K_{0}}^{\top}\bm{\Sigma}_{K_{0}}^{-1}\bm{a}_{K_{0}}>0$ and $\log N/N\to 0$. Thus $\mathbb{P}(\widehat{K}=k)\to 0$ for $k=1,\ldots,K_{0}-1$, and
$\lim_{N\to\infty}\mathbb{P}\left(\widehat{K}\geq K_{0}\right)=1-\lim_{N\to\infty}\sum_{k=1}^{K_{0}-1}\mathbb{P}\left(\widehat{K}=k\right)=1.$
@@
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x\right)$ $=\sum_{k=1}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\leq x,\widehat{K}=k\right)$
$\leq\sum_{k=1}^{K_{0}-1}\mathbb{P}\left(\widehat{K}=k\right)+\sum_{k=K_{0}}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\leq x\right)\to 0,$
since $\widehat{\Psi}^{2}_{k}\overset{p}{\to}\bm{a}_{k}^{\top}\bm{\Sigma}_{k}^{-1}\bm{a}_{k}>0$ for $k=K_{0},\ldots,D$. Therefore, for any $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathbb{P}\left(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x\right)=0,$
@@
A.8 Intuition behind the revised limiting null distribution (4.3)
We primarily follow the arguments in kallenberg1999data, janic2000data, and kraus2007data, and provide a heuristic derivation of the approximation $H(x)$ for the first case, as the latter two cases can be derived analogously. Note that
$\mathbb{P}\left(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x\right)=$ $\mathbb{P}\left(N\widehat{\Psi}_{1}^{2}\leq x,\widehat{K}=1\right)+\mathbb{P}\left(N\widehat{\Psi}_{2}^{2}\leq x,\widehat{K}=2\right)$
$+\mathbb{P}\left(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x,\widehat{K}\geq 3\right).$ (A.30)
The third term on the right-hand side of (A.8), $\mathbb{P}(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x,\widehat{K}\geq 3)$, can be neglected under $H_{0}$ (kallenberg1995data). The event $\{\widehat{K}=1\}$ is approximated by $\{N\widehat{\Psi}^{2}_{1}-\log N\geq N\widehat{\Psi}^{2}_{2}-2\log N\}=\{N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\widehat{K}=2\}$ is approximated by $\{N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})>\log N\}$.
We first investigate the asymptotic distribution of $(N\widehat{\Psi}^{2}_{1},N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1}))^{\top}$. The limiting distributions of $N\widehat{\Psi}^{2}_{1},N\widehat{\Psi}^{2}_{2}$ are functions of a bivariate normal vector $(R_{1},R_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\bm{\Sigma}_{2})$. We generally denote elements of $\bm{\Sigma}_{2}$ as $\begin{pmatrix}a&b\\ b&c\end{pmatrix}$ and $\rho=b/\sqrt{ac}$. The distribution of $(R_{1},R_{2})^{\top}$ can be constructed through two independent standard normal variables $G_{1},G_{2}$ if we denote $\widetilde{R}_{1}=\sqrt{a}(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})$ and $\widetilde{R}_{2}=\sqrt{c}G_{2}$, then $(\widetilde{R}_{1},\widetilde{R}_{2})^{\top}\overset{d}{=}(R_{1},R_{2})^{\top}$. Thus, $N\widehat{\Psi}^{2}_{1}\overset{d}{\to}R_{1}^{2}/a\overset{d}{=}\widetilde{R}_{1}^{2}/a=(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})^{2}$. For $N\widehat{\Psi}_{2}^{2}$, straightforward computations yield that
$N\widehat{\Psi}_{2}^{2}$ $\overset{d}{\to}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)\bm{\Sigma}_{2}^{-1}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(c\widetilde{R}_{1}^{2}-2b\widetilde{R}_{1}\widetilde{R}_{2}+a\widetilde{R}_{2}^{2}\right)$
@@
Then $N(\widehat{\Psi}_{2}^{2}-\widehat{\Psi}_{1}^{2})\overset{d}{\to}(\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{2}$. Since $(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2},\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\mathbf{I}_{2})$, we obtain that $N\widehat{\Psi}^{2}_{1}$, $N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})$ are both asymptotically $\chi_{1}^{2}$ distributed and are asymptotically independent.
@@
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,\widehat{K}=2\right)\approx\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}>\log N\right)=0,$
by $N\widehat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{1}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\leq\log N\right)$